OpenStudy (anonymous):

what is the quotient of 3/4 divided by 5/6

OpenStudy (anonymous):

Thtas 3/4 x 6/5

OpenStudy (ghostgate):

0.9 - You may either divide by turning the fractions into numbers which is done easily: 3/4 = 0.75 because divide 1 by 4 and multiply by 3 and you get 0.75 5/6 = 0.8333.. because divide 1 by 6 and multiply by 5 and you get 0.8333.. Take these numbers so you have 0.75 divided by 0.8333.. and your answer is 0.9 Or you may do what Easyaspi314 did which is: 3/4 x 6/5 which still equals 0.9 and has the same concept. Since we already know that 3/4 = 0.75 we just have to figure out what 6/5 is It equals 1.2 because divide 1 by 5 and multiply by 6 to get 1.2 so the answer is 0.75 x 1.2 which again equals 1.2 Hope this helps, G'day.

OpenStudy (ghostgate):

equals 0.9 sorry.

OpenStudy (texaschic101):

or just leave it as a fraction... 3/4 x 6/5 = 18/20 = 9/10

OpenStudy (ghostgate):

That too.